OpenStudy (anonymous):

Given the linear equation y=-3/5x-4 , find the y-coordinates of the points (-5, ), (0, ), and (5, ). Please show all of your work.

OpenStudy (anonymous):

for the point (-5, ) all you need to do is plug in a -5 for the x in the equation: y = -3/5*(-5) - 4 y = 3 - 4 y = -1 that's all you need to do for the others...

OpenStudy (anonymous):

You'd have to know the equation. Like if the linear equation was y = 1/5 x + 3, you'd put in -5 for x y = 1/5 • -5 + 3 y = -1 + 3 y = 2 so 2 goes with the -5 Then put in 0 for x: y = 1/5 • 0 + 3 which is 3 then y = 1/5 • 5 + 3 which is 1 + 3 = 4